OpenStudy (anonymous):

what are the solutions of 16t^2-5=4 i need to solve by using factoring the the 0 products

OpenStudy (anonymous):

16t^2-9=0 difference of 2 squares a^2-b^2=(a-b)(a+b) (4t-3)(4t+3)=0 principle of 0 products says that for the above equation to be true, 4t-3=0 or 4t+3=0